OpenStudy (anonymous):

Remmi wrote the equation of the line y=1/3(x+2).He solved for x and got x= 3y-2. Which of the following is an equivalent equation for x?

OpenStudy (lucaz):

what are the options?

OpenStudy (anonymous):

1. x= y- 11/3 2. x= y+ 7/3 3. x= 3(y- 2/3) 4. x= 3(y+ 2/3)

OpenStudy (lucaz):

3. x=3(y-2/3)

OpenStudy (anonymous):

thank you.